OpenStudy (anonymous):

Line A is represented by the equation given below: x + y = 4 What is most likely the equation for line B, so that the set of equations has infinitely many solutions? 4x + 4y = 4 4x + y = 4 2x + 2y = 8 x + y = 8

OpenStudy (hybrik):

To have infinite solutions, the slope and y intercept have to be the same

OpenStudy (anonymous):

a?

OpenStudy (hybrik):

The simplified way is y=-x+4

OpenStudy (anonymous):

so its a or b

OpenStudy (hybrik):

A

OpenStudy (hybrik):

Medal for helping?

OpenStudy (anonymous):

The work of a student to solve a set of equations is shown: Equation 1: m = 8 + 2n Equation 2: 3m = 4 + 4n Step 1: −3(m) = −3(8 + 2n) [Equation 1 is multiplied by −3.] 3m = 4 + 4n [Equation 2] Step 2: −3m = −24 − 6n [Equation 1 in Step 1 is simplified.] 3m = 4 + 4n [Equation 2] Step 3: −3m + 3m = −24 − 6n + 4n [Equations in Step 2 are added.] Step 4: 0 = −24 − 2n Step 5: n = −12 In which step did the student first make an error? Step 4 Step 3 Step 2 Step 1

OpenStudy (hybrik):

Step 3

OpenStudy (anonymous):

Variable g is 8 more than variable w. Variable g is also 4 less than w. Which pair of equations best models the relationship between g and w? g = w − 8 g = w + 4 w = 8g w = g − 4 g = w + 8 g = w − 4 w = 8g w = g + 4

OpenStudy (hybrik):

g = w + 8 g = w − 4